# Sample, Don’t Search: Rethinking Test-Time Alignment for Language Models

### v1

@@ Abstract
Increasing test-time computation has emerged as a promising direction for improving language model performance, particularly in scenarios where model finetuning is impractical or impossible due to computational constraints or private model weights. However, existing test-time search methods using a reward model (RM) often degrade in quality as compute scales, due to the over-optimization of what are inherently imperfect reward proxies. We introduce QAlign, a new test-time alignment approach. As we scale test-time compute, QAlign converges to sampling from the optimal aligned distribution for each individual prompt. By adopting recent advances in Markov chain Monte Carlo for text generation, our method enables better-aligned outputs without modifying the underlying model or even requiring logit access. We demonstrate the effectiveness of QAlign on mathematical reasoning benchmarks (GSM8K and GSM-Symbolic) using a task-specific RM, showing consistent improvements over existing test-time compute methods like best-of-$n$ and majority voting. Furthermore, when applied with more realistic RMs trained on the Tülu 3 preference dataset, QAlign outperforms direct preference optimization (DPO), best-of-$n$, majority voting, and weighted majority voting on a diverse range of datasets (GSM8K, MATH500, IFEval, MMLU-Redux, and TruthfulQA). A practical solution to aligning language models at test time using additional computation without degradation, our approach expands the limits of the capability that can be obtained from off-the-shelf language models without further training.
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable capabilities through learning from human preferences (Stiennon et al., 2022; Fernandes et al., 2023; Kaufmann et al., 2023; Peters & Schaal, 2007; Peng et al., 2019; Korbak et al., 2022b; a; Go et al., 2023). However, there are a number of problems with deploying a single aligned model: alignment approaches typically average multiple human preferences, constructing a monolithic preference model and target policy. Furthermore, approaches to adapting these models through finetuning have become increasingly impractical, requiring enormous computational resources. They are entirely impossible when model weights are private, as with many state-of-the-art models (OpenAI, 2024; Anthropic, 2024; Gemini, 2024).
While conventional alignment methods optimize a single model for aggregate performance across a distribution of prompts and then produce a random generation during inference, researchers have found great improvement from scaling the amount of compute expended at test time for each prompt (Brown et al., 2024; Snell et al., 2024). Here, multiple outputs are generated and then used to produce a final answer either via reward maximization (known as “best-of-$n$,” or BoN; Gao et al., 2022; Stiennon et al., 2020), majority voting (MV; Wang et al., 2023b) or weighted majority voting (WMV; Li et al., 2023). While promising, existing search-based methods face fundamental limitations (Liu et al., 2024; Wu et al., 2024; Zhang et al., 2023; Xie et al., 2023): as inference compute scales, these methods over-optimize learned reward models (RMs), which are inherently imperfect proxies (Gao et al., 2022).
In this paper, we introduce QAlign, a test-time alignment method that converges to sampling from the optimal aligned distribution (defined in §2 as the true target for existing RLHF methods) for each prompt, individually, as the test-time compute budget is increased. This enables more aligned responses without requiring access to the logits or training any underlying LM (only sampling from it). Specifically, we adapt Quest (Faria et al., 2024). Quest was designed to repurpose a language model, creating a Markov chain converging toward a (Gibbs) distribution defined by by machine translation quality estimation metrics. In this work, we show that we can use Quest to align LMs at test time according to an RM learned from preference data (in place of the quality estimation metric) and to provide a single good final prediction.
@@
In our experiments (§4), we compare QAlign with existing test-time compute methods on mathematical reasoning benchmarks (GSM8K, Cobbe et al., 2021, and GSM-Symbolic, Mirzadeh et al., 2024), by training a task-specific RM (§4.1), and using it to increase the capability of Llama3.1-8B-Instruct (Llama, 2024). We obtain a consistent reduction in error as we scale inference time computation across scales, as shown in Figure 2. Unlike BoN, QAlign’s performance does not degrade as we increase the compute per question. Furthermore, when applied with more realistic RMs (§ 4.2), trained on the Tülu3 (Lambert et al., 2024) preference dataset, where over-optimization starts to appear at a lower inference budget, and applied to Tülu3-8B-SFT, our experiments show that QAlign can consistently outperform direct preference optimization (DPO; Rafailov et al., 2024), which updates the model weights, as well as BoN, MV, and WMV on a diverse range of datasets in mathematical reasoning (GSM8K and MATH500;Hendrycks et al., 2021; Lightman et al., 2023), knowledge recall (MMLU-Redux; Gema et al., 2025), TruthfulQA (Lin et al., 2022) and instruction following (IFEval; Zhou et al., 2023b), as shown in Figure 1.111www.questdecoding.com/alignment
@@
  • We show a $35\%$ increase in accuracy on GSM8K and $98\%$ on GSM-Symbolic relative to a single generation, using a task-specific RM, beating BoN, MV, and WMV.
  • We show that QAlign outperforms DPO when applied to Tülu3 family of models, in general RMs trained on the Tülu3 preference dataset, even when compared with the same compute budget at test-time via MV, achieving a $57\%$ increase in average accuracy relative to a single generation across a suite of benchmarks.
@@ 2 Background: Language Model Alignment
Generating text from autoregressive LMs involves conditioning on a prompt $x$ encoding the particular question we want to solve. Let $\mathcal{Y}$ denote the set of possible LM responses (outputs). The distribution over responses $y\in\mathcal{Y}$, given the prompt $x$, can be factorized as the product of conditional probabilities over individual tokens $\langle y_{1},y_{2},\ldots,y_{N}\rangle$, where each $y_{i}\in\mathcal{V}$, a fixed vocabulary. The probability of a particular response $y$ for a given input $x$ can be written as
$p_{\text{LM}}(y\mid x)=\textstyle\prod_{i=1}^{N}p_{\text{LM}}(y_{i}\mid y_{<i},x).$ (1)
Here, $y_{<i}=\langle y_{1},y_{2},\ldots,y_{i-1}\rangle$.
Finetuning a pretrained LM $p_{\text{LM}}(y\mid x)$ to align with preferences encoded by a reward function $r(y,x)$ can be cast as Bayesian inference (Korbak et al., 2022a). We seek $\pi(y\mid x)$, which is initially equal to a prior $p_{\text{LM}}(y\mid x)$, and updated to conform to evidence provided by the human preferences model, where $p(\gamma=1\mid y,x)\propto\exp(r(y,x)/\beta)$ is the likelihood that the prompt-response pair $(x,y)$ is human-preferred. Using Bayes’ rule, the posterior takes the following form:
@@
To approximate $\pi(y\mid x)$, many methods have been proposed (Stiennon et al., 2022; Fernandes et al., 2023; Kaufmann et al., 2023; Peters & Schaal, 2007; Peng et al., 2019; Korbak et al., 2022b; a; Go et al., 2023). Given a dataset of prompts $\mathcal{D}$, one common approach is to estimate a parameterized variational approximation $q_{\theta}(y\mid x)$:
@@
Methods like PPO (Schulman et al., 2017) and others (Shao et al., 2024; Rafailov et al., 2024; Ahmadian et al., 2024) optimize this objective with low-variance gradient estimates, resulting in stable optimization. Importantly, $q_{\theta}(y\mid x)$ is a newly trained LM initialized with $p_{\text{LM}}(y\mid x)$. The large-scale efforts to align language models generally alternate between learning $q_{\theta}(y\mid x)$ and $r(y,x)$, with human preferences on samples generated from $q_{\theta}(y\mid x)$ (Llama, 2024; 2023; Gemini, 2024). This avoids allowing the approximate posterior to exploit regions where the reward is out of distribution and also to refine the resolution of the reward towards promising regions.
@@ 3 Test-Time Alignment via MCMC
Casting language model alignment as posterior inference decouples our target goal from the procedure used to achieve it. While current approaches rely on variational inference to learn a single model $q_{\theta}(y\mid x)$ (Eq. 3), this strategy faces four limitations. First, it requires expensive model finetuning, which is costly with large models. Second, many models’ weights are not openly shared, including those of state-of-the-art models like GPT-4 (OpenAI, 2024) and Gemini (Gemini, 2024). Third, amortizing the approximation across all prompts necessarily sacrifices the quality of approximation for any individual prompt $x$ to achieve good average performance. Fourth, the approach assumes a monolithic notion of human preferences encoded in $r(y,x)$, offering no flexibility to adapt to varying user preferences or contexts at inference time.
These limitations motivate a shift toward local posterior approximations, where we achieve a better approximation as we increase the compute budget at test time on a single prompt $x$. As we will describe in (§3.1), our key insight is that rather than estimating a single parametric approximation, we can use recent advancements in Markov chain Monte Carlo (MCMC) sampling with LMs (Faria et al., 2024) and obtain a sequence of samples $\mathcal{S}=\langle y^{0},y^{1},\dots,y^{T}\rangle$ from $\pi(y\mid x)$ and use them for optimal decision-making.
A principled way to select a final response from the generated candidate set $\mathcal{S}$ from $\pi(y\mid x)$ is by selecting the most common element (mode) i.e., via majority voting (MV; Wang et al., 2023b). However, this only works for tasks with well-defined answers, such as mathematical reasoning or multiple-choice questions. For open-ended generation tasks, we apply a generalization of this approach through the minimum Bayes risk (MBR) principle (Kumar & Byrne, 2002; Eikema & Aziz, 2020; Farinhas et al., 2023; Bertsch et al., 2023). The MBR framework selects the output $\hat{y}$ that maximizes the expected task-specific utility $u(y,y^{\prime})$:
@@
When the utility metric $u(y,y^{\prime})$ is defined as an exact match over the final answer, MBR amounts to a MV strategy over $\mathcal{S}$. We employ ROUGE (Lin, 2004) as our utility metric for open-ended generation tasks, following Bertsch et al. (2023). While this results in $O(T^{2})$ similarity computations, ROUGE is computationally lightweight, and efficient low-rank approximations (Trabelsi et al., 2024) can be used when dealing with heavier neural-based metrics.
@@ 3.1 MCMC for Text Generation
With the goal of generating a sequence of samples from $\pi_{\beta}(y\mid x)$, we will construct a Markov chain $(y^{0},y^{1},\ldots,y^{T})$ that has $\pi_{\beta}(y\mid x)$ as its equilibrium distribution. The chain starts from a hypothesis $y^{0}\sim p_{\text{LM}}(y\mid x)$. On the $t$th iteration, it draws a new hypothesis $y$ from a proposal distribution $q(y\mid y^{t},x)$ , and this hypothesis is accepted with an acceptance probability $\alpha_{\beta}(y,y^{t})\in[0,1]$. The proposal distribution $q(y\mid y^{t},x)$ we use is the one proposed by Quest (Faria et al., 2024). It is based on a LM that samples suffixes $p_{\text{LM}}(y_{i:N}\mid y_{<i},x)$ starting at uniformly sampled index $i$. This proposal can be written as
$q(y\mid y^{t},x,i)=p_{\text{LM}}(y_{i:N}\mid y^{t}_{<i},x)\times\boldsymbol{1}\{y_{1:j}=y_{1:j}^{t}\},$ (5)
Following the Metropolis-Hastings algorithm (MH; Hastings, 1970), the acceptance probability is
@@
MH converges to the unique stationary distribution $\pi_{\beta}(y\mid x)$, regardless of the initial distribution, because the transition distribution of the Markov chain, $p(y^{t}\mid y^{t-1},x)$ which results from generating a candidate from $q(y^{t}\mid y^{t-1},x)$ followed by an accept/reject step, satisfies the Markov chain ergodic theorem (Neal, 2011).
@@
While more complex proposals could be built based on LMs prompted to self-refine (Zhou et al., 2023a; Madaan et al., 2023; Yao et al., 2023), recent work demonstrates that even strong LMs often fail to improve reasoning due to models’ inability to gauge the correctness of their outputs and localize errors (Huang et al., 2024). Because of this, a growing body of research tries to teach LMs how to self-correct by creating data-augmentation recipes (Welleck et al., 2022; Havrilla et al., 2024) or new datasets (Wang et al., 2023a; Chen et al., 2024; Lee et al., 2024; Saunders et al., 2022; Schick et al., 2022). However, when the proposal is different from the base model, we lose the simplicity of the acceptance criterion in Eq. 9 and need to keep track of two sets of logits to calculate the acceptance criterion in Eq. 6.
As our approach only requires access to the ability to generate continuations given a prefix (not access to base model parameters or logits), it can also be used with closed LMs through an API. However, we limit our experiments to open-weight models. We provide in Appendix B an analysis of the computational cost associated with running QAlign.
@@ 3.2 Importance Sampling
A natural competing approach to QAlign for approximating the intractable expectation in Eq. 4 is to use importance sampling (IS; Kahn, 1950). Rather than attempting to directly sample from the target distribution $\pi(y\mid x)$, IS generates samples from the base LM $p_{\text{LM}}(y\mid x)$ and reweights them to match the target distribution $\pi(y\mid x)$:
$\mathbb{E}_{y\sim\pi(y\mid x)}[u(y,y^{\prime})]$ $=\mathbb{E}_{y\sim p_{\text{LM}}(y\mid x)}\left[\frac{\pi(y\mid x)}{p_{\text{LM}}(y\mid x)}u(y,y^{\prime})\right]=\mathbb{E}_{y\sim p_{\text{LM}}(y\mid x)}\left[\frac{\exp(\frac{1}{\beta}r(y,x))}{Z_{\beta}(x)}u(y,y^{\prime})\right]$
Note that the partition function $Z_{\beta}(x)$ can be written as an expectation over samples from the base model. So in the end, making this approximation boils down to the following steps:
@@
In the LM literature, this procedure, for the specific case of tasks with well-defined final answers (i.e., the utility metric $u(y,y^{\prime})$ is defined as an exact match), the MBR output is the same as weighted majority voting (WMV; Li et al., 2023). Similar to QAlign, WMV is guaranteed to converge to the optimal decision rule as computational resources increase. However, because it relies solely on independent samples drawn from the base LM, the generation process cannot be directly steered toward more promising regions of the output space. This limitation becomes particularly problematic when the base LM infrequently produces high-reward responses. In such cases, the approximation of the target expectation, and corresponding final decision, will be compromised.

### v2

@@ Abstract
Increasing test-time computation has emerged as a promising direction for improving language model performance, particularly in scenarios where model finetuning is impractical or impossible due to computational constraints or private model weights. However, existing test-time search methods using a reward model (RM) often degrade in quality as compute scales, due to the over-optimization of what are inherently imperfect reward proxies. We introduce QAlign, a new test-time alignment approach. As we scale test-time compute, QAlign converges to sampling from the optimal aligned distribution for each prompt. By adopting recent advances in Markov chain Monte Carlo for text generation, our method enables better-aligned outputs without modifying the underlying model or even requiring logit access. We demonstrate the effectiveness of QAlign on mathematical reasoning benchmarks (GSM8K and GSM-Symbolic) using a task-specific RM, showing consistent improvements over existing test-time compute methods like best-of-$n$ and majority voting. When applied with more realistic RMs trained on the Tülu 3 preference dataset, QAlign outperforms direct preference optimization (DPO), best-of-$n$, majority voting, and weighted majority voting on a diverse range of datasets (GSM8K, MATH500, IFEval, MMLU-Redux, and TruthfulQA). A practical solution to aligning language models at test time using additional computation without degradation, our approach expands the limits of the capability that can be obtained from off-the-shelf language models without further training.
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable capabilities through learning from human preferences (firstrlhf; fernandes2023bridging; kaufmann2023survey; 10.1145/1273496.1273590; peng2019advantageweighted; rlhfandbayes; korbak2022reinforcement; go2023aligning). However, there are a number of problems with deploying a single aligned model: alignment approaches typically average multiple human preferences, constructing a monolithic preference model and target policy. Furthermore, approaches to adapting these models through finetuning have become increasingly impractical, requiring enormous computational resources. They are entirely impossible when model weights are private, as with many state-of-the-art models (openai2024gpt4technicalreport; TheC3; geminiteam2024geminifamilyhighlycapable).
While conventional alignment methods optimize a single model for aggregate performance across a distribution of prompts and then produce a random generation during inference, researchers have found great improvement from scaling the amount of compute expended at test time for each prompt (brown2024largelanguagemonkeysscaling; snell2024scalingllmtesttimecompute). Here, multiple outputs are generated and then used to produce a final answer either via reward maximization (known as “best-of-nn,” or BoN; gao2022scalinglawsrewardmodel; stienon2020learning; fern2022qualityaware), majority voting (MV; selfconsistency) or weighted majority voting (WMV; li-etal-2023-making). While promising, existing search-based methods face fundamental limitations (liu2024dontthrowawayvalue; wu2024inferencescalinglawsempirical; zhang2023planninglargelanguagemodels; xie2023selfevaluationguidedbeamsearch): as inference compute scales, these methods over-optimize learned reward models (RMs), which are inherently imperfect proxies (gao2022scalinglawsrewardmodel).
In this paper, we introduce QAlign, a test-time alignment method that converges to sampling from the optimal aligned distribution (defined in §1 as the true target for existing RLHF methods) for each prompt, individually, as the test-time compute budget is increased. This enables more aligned responses without requiring access to the logits or training any underlying LM (only sampling from it). Specifically, we adapt Quest (quest). Quest was designed to repurpose a language model, creating a Markov chain converging toward a (Gibbs) distribution defined by machine translation quality estimation metrics. In this work, we show that we can use Quest to align LMs at test time according to an RM learned from preference data (in place of the quality estimation metric) and to provide a single good final prediction.
@@
In our experiments (§4), we compare QAlign with existing test-time compute methods on mathematical reasoning benchmarks (GSM8K, gsm8k, and GSM-Symbolic, gsmsymb), by training a task-specific RM (§4.1), and using it to increase the capability of Llama3.1-8B-Instruct (llama3). We obtain a consistent reduction in error as we scale inference time computation across scales, as shown in Figure 2. Unlike BoN, QAlign’s performance does not degrade as we increase the compute per question. Furthermore, when applied with more realistic RMs (§ 4.2), trained on the Tülu3 (tulu3) preference dataset, where over-optimization starts to appear at a lower inference budget, and applied to Tülu3-8B-SFT, our experiments show that QAlign can consistently outperform direct preference optimization (DPO; dpo), which updates the model weights, as well as BoN, MV, and WMV on a diverse range of datasets in mathematical reasoning (GSM8K and MATH500; hendrycksmath2021; lightman2023letsverifystepstep), knowledge recall (MMLU-Redux; gema2025mmlu), TruthfulQA (truthfulqa) and instruction following (IFEval; ifeval), as shown in Figure 1.111Website with code and additional resources will be provided upon acceptance.
@@
  • We show a $28.4\%$ increase in accuracy on GSM8K and $78.34\%$ on GSM-Symbolic relative to a single generation, using a task-specific RM, beating BoN, MV, and WMV.
  • We show that QAlign outperforms DPO when applied to Tülu3 family of models, in general RMs trained on the Tülu3 preference dataset, even when compared with the same compute budget at test-time via MV, achieving a $47.85\%$ increase in average accuracy relative to a single generation across a suite of benchmarks.
@@ 2 Background: Language Model Alignment
Finetuning a pretrained LM $p_{\text{LM}}(y\mid x)$ to align with preferences encoded by a reward function $r(y,x)$ can be cast as Bayesian inference (korbak2022reinforcement). We seek $\pi(y\mid x)$, which is initially equal to a prior $p_{\text{LM}}(y\mid x)$, and updated to conform to evidence provided by the human preferences model. Let $\gamma=1$ denote the event that response $y$ is maximally preferred for prompt $x$ (i.e., $y\succeq y^{\prime}$ for all $y^{\prime}\in\mathcal{Y}$). Assuming the reward function is bounded, we can model the likelihood of this event as $p(\gamma=1\mid y,x)=\exp\left(\left({r(y,x)-\max_{y^{\prime}}r(y^{\prime},x)}\right)/{\beta}\right)$. Using Bayes’ rule, the posterior takes the following form:
@@
To approximate $\pi(y\mid x)$, many methods have been proposed (firstrlhf; fernandes2023bridging; kaufmann2023survey; 10.1145/1273496.1273590; peng2019advantageweighted; rlhfandbayes; korbak2022reinforcement; go2023aligning). Given a dataset of prompts $\mathcal{D}$, one common approach is to estimate a parameterized variational approximation $q_{\theta}(y\mid x)$(see Appendix B for details):
@@
Methods like PPO (schulman2017proximalpolicyoptimizationalgorithms) and others (shao2024deepseekmathpushinglimitsmathematical; dpo; rloo) optimize this objective with low-variance gradient estimates, resulting in stable optimization. Importantly, $q_{\theta}(y\mid x)$ is a newly trained LM initialized with $p_{\text{LM}}(y\mid x)$.
@@ 3 Test-Time Alignment via MCMC
Casting language model alignment as posterior inference decouples our target goal from the procedure used to achieve it. While current approaches rely on variational inference to learn a single model $q_{\theta}(y\mid x)$ (Eq. 2), this strategy faces four limitations. First, it requires expensive model finetuning, which is costly with large models. Second, many models’ weights are not openly shared, including those of state-of-the-art models like GPT-4 (openai2024gpt4technicalreport) and Gemini (geminiteam2024geminifamilyhighlycapable). Third, amortizing the approximation across all prompts necessarily sacrifices the quality of approximation for any individual prompt $x$ to achieve good average performance. Fourth, the approach assumes a monolithic notion of human preferences encoded in $r(y,x)$, offering no flexibility to adapt to varying user preferences or contexts at inference time.
These limitations motivate a shift toward local posterior approximations, where we achieve a better approximation as we increase the compute budget at test time on a single prompt $x$. As we will describe in (§3.1), our key insight is that rather than estimating a single parametric approximation, we can use recent advancements in Markov chain Monte Carlo (MCMC) sampling with LMs (quest) and obtain a sequence of samples $\mathcal{S}=\langle y^{0},y^{1},\dots,y^{T}\rangle$ from $\pi(y\mid x)$ and use them for optimal decision-making.
A principled way to select a final response from the generated candidate set $\mathcal{S}$ from $\pi(y\mid x)$ is by selecting the most common element (mode) i.e., via majority voting (MV; selfconsistency). However, this only works for tasks with well-defined answers, such as mathematical reasoning or multiple-choice questions. For open-ended generation tasks, we apply a generalization of this approach through the minimum Bayes risk (MBR) principle (10.3115/1118693.1118712; eikema-aziz-2020-map; farinhas2023empiricalstudytranslationhypothesis; bertsch2023itsmbrwaydown). The MBR framework selects the output $\hat{y}$ that maximizes the expected task-specific utility $u(y,y^{\prime})$:
@@
When the utility metric $u(y,y^{\prime})$ is defined as an exact match over the final answer, MBR amounts to a MV strategy over $\mathcal{S}$. We employ ROUGE (lin-2004-rouge) as our utility metric for open-ended generation tasks, following bertsch2023itsmbrwaydown. While this results in $O(T^{2})$ similarity computations, ROUGE is computationally lightweight.
@@ 3.1 MCMC for Text Generation
With the goal of generating a sequence of samples from $\pi_{\beta}(y\mid x)$, we will construct a Markov chain $(y^{0},y^{1},\ldots,y^{T})$ that has $\pi_{\beta}(y\mid x)$ as its equilibrium distribution. The chain starts from a hypothesis $y^{0}\sim p_{\text{LM}}(y\mid x)$. On the $t$th iteration, it draws a new hypothesis $y$ from a proposal distribution $q(y\mid y^{t},x)$ , and this hypothesis is accepted with an acceptance probability $\alpha_{\beta}(y,y^{t})\in[0,1]$. The proposal distribution $q(y\mid y^{t},x)$ we use is the one proposed by Quest (quest). It is based on a LM that samples suffixes $p_{\text{LM}}(y_{i:N}\mid y_{<i},x)$ starting at uniformly sampled index $i$. This proposal can be written as
$q(y\mid y^{t},x,i)=p_{\text{LM}}(y_{i:N}\mid y^{t}_{<i},x)\times\boldsymbol{1}\{y_{1:i}=y_{1:i}^{t}\},$ (4)
Following the Metropolis-Hastings algorithm (MH; hastings), the acceptance probability is
@@
MH converges to the unique stationary distribution $\pi_{\beta}(y\mid x)$, regardless of the initial distribution, because the transition distribution of the Markov chain, $p(y^{t}\mid y^{t-1},x)$ which results from generating a candidate from $q(y^{t}\mid y^{t-1},x)$ followed by an accept/reject step, satisfies the Markov chain ergodic theorem (Neal2011ProbabilisticIU)(see Appendix C for a detailed proof).
@@
While more complex proposals could be built based on LMs prompted to self-refine (solvingchallengingmathword; selfrefine; react), recent work demonstrates that even strong LMs often fail to improve reasoning due to models’ inability to gauge the correctness of their outputs and localize errors (cantselfcorrect). Because of this, a growing body of research tries to teach LMs how to self-correct by creating data-augmentation recipes (welleckselfcorrect; glore) or new datasets (wang2023shepherdcriticlanguagemodel; chen2024improvingcodegenerationtraining; lee2024platypusquickcheappowerful; saunders2022selfcritiquingmodelsassistinghuman; schick2022peercollaborativelanguagemodel). However, when the proposal is different from the base model, we lose the simplicity of the acceptance criterion in Eq. 8 and need to keep track of two sets of logits to calculate the acceptance criterion in Eq. 5.
As our approach only requires access to the ability to generate continuations given a prefix (not access to base model parameters or logits), it can also be used with closed LMs through an API. However, we limit our experiments to open-weight models. We provide in Appendix D an analysis of the computational cost associated with running QAlign.
@@ 3.2 Importance Sampling
A natural competing approach to QAlign for approximating the intractable expectation in Eq. 3 is to use importance sampling (IS; is). Rather than attempting to directly sample from the target distribution $\pi(y\mid x)$, IS generates samples from the base LM $p_{\text{LM}}(y\mid x)$ and reweights them to match the target distribution $\pi(y\mid x)$:
$\mathbb{E}_{y\sim\pi(y\mid x)}[u(y,y^{\prime})]$ $=\mathbb{E}_{y\sim p_{\text{LM}}(y\mid x)}\left[\frac{\pi(y\mid x)}{p_{\text{LM}}(y\mid x)}u(y,y^{\prime})\right]=\mathbb{E}_{y\sim p_{\text{LM}}(y\mid x)}\left[\frac{\exp(\frac{1}{\beta}r(y,x))}{Z_{\beta}(x)}u(y,y^{\prime})\right].$
The partition function $Z_{\beta}(x)$ can be reframed as an expectation over samples from the base model: $Z_{\beta}(x)=\sum_{y\in\mathcal{Y}}p_{\text{LM}}(y\mid x)\exp(r(y,x)/\beta)=\mathbb{E}_{y\sim p_{\text{LM}}(y\mid x)}[\exp(r(y,x)/\beta)]$. This motivates a self-normalizing importance sampling approach, where we approximate both the numerator and denominator using samples $\{y^{(i)}\}_{i=1}^{K}\sim p_{\text{LM}}(y\mid x)$. The self-normalized estimator is consistent, but introduces bias due to the correlation between the numerator and denominator (owen2013monte). In the end, making this approximation boils down to the following steps:
@@
In the LM literature, this procedure, for the specific case of tasks with well-defined final answers (i.e., the utility metric $u(y,y^{\prime})$ is defined as an exact match), the MBR output is the same as weighted majority voting (WMV; li-etal-2023-making). Similar to QAlign, WMV is guaranteed to converge to the optimal decision rule as computational resources increase. However, because it relies solely on independent samples drawn from the base LM, the generation process cannot be directly steered toward more promising regions of the output space. This limitation becomes particularly problematic when the base LM infrequently produces high-reward responses. In such cases, the approximation of the target expectation, and corresponding final decision, will be compromised.